The frequency of the middle E note on a piano is 329.63 Hz. What is the wavelength of this note in centimeters? The speed of sound in air is 343.06 m/s.
1
Expert's answer
2020-10-27T12:41:32-0400

Q140734

The frequency of the middle E note on a piano is 329.63 Hz. What is the wavelength of this note in centimeters? The speed of sound in air is 343.06 m/s.


Solution: 


In question, we are given the frequency of the middle E note and we are also provided the speed of the sound. We can use the formula which relates the frequency and wavelength for finding the wavelength of the middle E note. 


The formula is 


 v = f * λ

  

where v is the speed of sound (m/s)

 f is the frequency (Hz or s-1 ) 

 and λ is the wavelength ( m ) 


Step 1: Find the wavelength of the note. 


Substitute, v = 343.06m/s and f = 329.63Hz in the formula we have 


343.06m/s = 329.63 s-1 * λ 


divide both the side by 329.63 s-1  


343.06m/s/ 329.63 s-1  =  329.63 s-1 * λ / 329.63 s-1


1.0407 m = λ ; 


Hence, wavelength (λ) = 1.0407 m. 


Step 2: Convert the wavelength to ‘centimeters’ 


In question, we are asked to find the wavelength in centimeters. 


We can use the conversion factor, 1m = 100cm for converting the wavelength to centimeters. 


Wavelength in ‘cm’ = 1.0407 m *  100cm/1m = 104.07 cm 


In question, we are given 329.63 Hz in 5 significant figure, so our final answer must also be in 5 significant figure.


104.07 cm is already in 5 significant figure.


 Hence the frequency of the middle E note in centimeters is 104.07 cm